[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sun May 07, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sun May 7 19:53:40 UTC 2017


HSU Shift Log for Sun May 07 2017 15:53:37 GMT-0400 (EDT)

Shift: H

Forecaster: J Beven

Technical Issues Summary:

ATCF issues: None.

N-AWIPS issues: None.

AWIPS-2/NCP issues: None.

Other Techincal issues: Updated: Two-day graphic TWO problem fixed. Issue  
was operator error during the creation of the graphical TWO. :-)

Miscellaneous issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170507/63266712/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list